### v3

# Pragmatic Inference for Moral Reasoning Acquisition: Generalization via Distributional Semantics

Abstract
Moral reasoning has emerged as a promising research direction for Large Language Models (LLMs), yet achieving generalization remains a central challenge. From a linguistic standpoint, this difficulty arises because LLMs are adept at capturing distributional semantics, which fundamentally differs from the morals which operate at the pragmatic level. This paper investigates how LLMs can achieve generalized moral reasoning despite their reliance on distributional semantics. We propose pragmatic inference methods grounded in moral foundations theory, which leverage contextual information at each step to bridge the pragmatic gap and guide LLMs in connecting moral foundations with moral reasoning objectives. Experimental results demonstrate that our approach significantly enhances LLMs’ generalization in moral reasoning, providing a foundation for future moral reasoning research based on pragmatic inference. Warning: this paper contains offensive language.
1 Introduction
With the advancement of Large Language Models (LLMs), how to enhance their capability of moral reasoning in order to regulate their performance in accordance with human moral values has attracted scholarly interest. However, it still remains as a challenging task due to the gap between distributional semantics used by LLMs and the pragmatics-based nature of moral reasoning  Liu et al. (2025b); metaphorically speaking, the gap between what is said and what is morally implied. Distributional semantics follow Firth’s (1957) distributional hypothesis which argues that linguistic units have similar meaning if they occur in similar contexts. In contrast, moral reasoning goes beyond this structuralism view of language and requires the capability of deriving conclusions from implicature of language use, moral norms and behavioral consequences, and observations of discursive rules of conduct Spencer-Oatey and Xing (2019). These unsaid variables need to be linguistically manifested so that the gap between distributional semantics and pragmatics in moral reasoning can be closed, which increases the methodological generalizability that previous studies have thus far struggled to achieve Rao et al. (2023); Preniqi et al. (2024); Jiang et al. (2025).
Prompt: Do people actually shit themselves when
they get very scared?
Reply: I don’t think I’ve ever heard of that
happening, but I guess it’s possible.
Rule of Thumb: It is wrong to shit on ourselves.
Moral Judgment: Agree.
Moral Foundation: care.
Table 1: Moral Reasoning Example From Ziems et al. (2022). The Moral Judgment for the given reply is Agree, and the Rule of Thumb (RoT) represents a judgment regarding an action mentioned in the prompt-reply, grounded in the moral foundation of fairness.
Table 1 demonstrates an example of the several elements that are involved in moral reasoning: rule of thumb (RoT), moral foundation, and moral judgment Ziems et al. (2022). Thus far, learning paradigms, which aim to equip LLMs with moral reasoning, generally train models to directly map a situation (Prompt and Reply) to moral judgment, with or without a reference to moral foundations and/or RoT  Hendrycks et al. (2020); Forbes et al. (2020); Jiang et al. (2021); Ziems et al. (2022); Jiang et al. (2025). Direct mapping often requires a huge amount of data to enable LLMs to capture the statistical correlations between situations and moral reasoning objectives.
On the other hand, we have witnessed the effectiveness of pragmatic inference used in other morality-relevant tasks, such as the identification of social biases and toxicity  Sap et al. (2020); Zhou et al. (2023); Chen and Wang (2025). For example,  Sap et al. (2020) developed a structured inference path that enables the identification of stereotypical implications encoded in explicit toxic language. Chen and Wang (2025) developed an inference chain that provides LLMs with the freedom to interpret online neologisms and rephrase the literal meanings of an input. Their method improved the performance of LLMs in recognizing highly implicit toxic language by 12% to 20%. These approaches are consonant with our goal of bridging the gap between distributional semantics and pragmatics, for instance, through interpreting literal meanings Chen and Wang (2025).
However, moral reasoning requires different inferential methods from the aforementioned toxicity or bias identification. Unlike toxic language (e.g., “lowering employment standards for females” implying “females are less capable”), the moral foundations***Please refer to Section 2 for more discussion on Moral Foundations Theory for moral reasoning. that underlie a moral situation are often not encoded by certain linguistic cues at all  Clark and Schaefer (1989); Graesser et al. (1994); Sap et al. (2022). Inferring them necessitates context-specific interpretations of abstract moral rules (e.g., the rule “do not hurt others” may manifest as “do not push another child to the floor” on a children’s playground and “do not emotionally abuse your partner” for adult relationships). Therefore, abstract moral foundations, rules of conduct, and context all need to be sufficiently textualized so as to be incorporated in machines’ reasoning process.
In this paper, we address the pragmatic gaps in moral reasoning by first proposing two pragmatic inference methods: (1) for classifying moral foundations (MFs) from rule of thumbs (RoTs), and (2) for making moral judgments based on a given situation and its underlying moral foundations. Building on this, we further introduce a third pragmatic inference method that allows LLMs to automatically infer moral foundations and then make moral judgments for any given situation. The three tasks move from extracting common principles from abstract rules (RoTs to MFs) to making judgment based on concrete situations. On the one hand, they approximate real-world scenarios where the inputs to moral reasoning may vary; on the other hand, they demonstrate the flexibility of these methods across different moral reasoning tasks, from identifying moral foundations to making judgments.
Experimental results show that our proposed pragmatic inference methods significantly improves generalization in moral reasoning tasks. Section 3 presents the benchmark annotation and experimental settings. Section 4 introduces our proposed pragmatic inference methods. Section 5 presents further analysis to characterize the achieved generalization with our proposed pragmatic inference methods. Finally, Section 7 concludes our findings and outlines potential directions for future work.
2 Related Works
Distributional Semantics is one of the theoretical foundations for embeddings. It interprets the meaning of a linguistic unit by additional signs (e.g., words, emoji) surrounding it Harris (1954); Firth (1957); McDonald and Ramscar (2001); Lenci and Sahlgren (2023); Boleda (2020). However, due to its nature as a structuralism approach to understanding language, distributional semantics suffers limitations. It treats the “context” purely as the texts provided, neglecting the many “unsaid” and latent parts of real interactions, such as the social constraints that a speaker encounters, the effects that a speech creates on the listeners, and the delivery of intention Lenci and others (2008); Purver and Sadrzadeh (2015). These are studied in pragmatics, where LLMs have shown several deficiencies, e.g., understanding implicature Cong (2024), comprehending metaphors and humours Ruis et al. (2023); Barattieri di SanPietro et al. (2023), and moral alignment Khamassi et al. (2024); Liu et al. (2024).
Pragmatic Inference and Moral Reasoning. In this work, we leverage pragmatic inference to facilitate LLMs to achieve stronger performance in moral reasoning. Pragmatic inference refers to the process of deriving conclusions of meanings based on context, intention, and language use Elder (2024). The inferential process is often not verbally realized and, instead, involves evaluations made at the metapragmatic level Kádár and Haugh (2013). That is, humans construct meta links that connect an instance of language use to certain social qualities (e.g., linking the phrase “thank you” to the concept of politeness). They learn these metapragmatic links over their physical and social experiences, which are entirely absent in LLMs. These limitations suggest that: (i) the inferential process, that is not verbally realized and often does not need to be for humans, needs to be textualized so that LLMs can learn; and (ii) the context, where humans gain their experiences, needs to be sufficiently textualized, including both societal (e.g., social norms) and situational contexts (e.g., situated interpretations of language use). Without effective methods to textualize them, it is not a surprise that moral reasoning has long been a challenge in the broader field of artificial intelligence Allen et al. (2005); Van Rysewyk and Pontier (2014); Wallach et al. (2020).
Moral Foundations Theory (MFT). Haidt and Joseph (2004) and Graham et al. (2013) describe a framework for understanding human morals and moral reasoning. MFT hypothesizes that moral judgments are guided by domain-specific foundations, i.e., care, fairness, loyalty, liberty, authority, and sanctity. These foundations capture the diversity of moral concerns and have been widely adopted in computational social science Johnson and Goldwasser (2018); Wendell and Tatalovich (2021); Reinig et al. (2024). They provide a principled way to annotate, analyze, and model moral dimensions in language, making them particularly relevant for studying how LLMs represent and reason about morality Forbes et al. (2020); Ziems et al. (2022).
3 Preliminary
Benchmark. In this paper, we adopt the Moral Integration Corpus (MIC) benchmark and its annotations from Ziems et al. (2022). We chose the MIC benchmark because it closely reflects how humans make moral judgments in daily communication and provides richer contextual information. One advantage of the MIC dataset is that it provides levels of agreement between the annotators, which allows us to focus on the less controversial part of their data. Table 1 presents an example from MIC. The example consists of a Prompt (“Do people actually shit themselves when they get very scared??”) and a Reply (“I don’t think I’ve ever heard of that happening, but I guess it’s possible..”), with three evaluations annotated: agreement with the reply as moral judgment, RoT, and MF. The MIC benchmark utilizes the six MFs defined by Graham et al. (2013): care, fairness, liberty, loyalty, authority, and sanctity. We take two of the evaluations as our objectives: moral judgment (agree, neutral, or disagree with the reply) and MFs. RoT, instead, is used as part of the learning paradigm. RoT is not selected as a learning objective because: (i) it does not have a fixed answer and, more crucially, multiple RoTs may exist in one situationWe refer readers to Section 3.1 of Ziems et al. (2022) for further details., and (ii) it has a rather dynamic relationship with MFs.
Experimental Settings. Following the steps of our pragmatic inference frameworks (detailed in Section 4), we design several questions and prompt the deepseek model to answer them. The step-wise answers are then used as the training dataThe data used in our experiments is available in the anonymized GitHub repository: https://anonymous.4open.science/r/moralpragmatics-F891. for fine-tuning the LLMs towards moral reasoning objectives. We chose Llama3-3.2-1B, Llama3-3.2-3B Grattafiori et al. (2024) and Mistral-7B (Mistral-7B-v0.3 Jiang et al. (2023)) as our backbone models, and leverage the supervised fine-tuning objective to train those models by following previous studies Ziems et al. (2022); Liu et al. (2025b). We intentionally select smaller models, as they are generally considered less capable than their larger counterparts. We aim to demonstrate the effectiveness of our pragmatic inference frameworks through the strong performance achieved by these smaller models.
To reiterate, we have three pragmatic inference frameworks, corresponding to three tasks. For the moral judgment task (Agree, Neutral, or Disagree with the reply), we use prediction accuracy as the evaluation metric. For the moral foundations classification (MFC) task, we compute prediction accuracy according to the number of MFs. Specifically, accuracies are reported separately for cases where a single RoT corresponds to one MF and for cases referring to two or more MFs. In other words, we have taken into consideration the one-to-many relationship between RoT/situation and MFs. We denote the LLM $f$ is parameterized by $\theta$, and the definitions (see Figure 5) of MFs are represented with $d_{m}$. Let $(x_{S},x_{R},y_{M},y_{J})$ denote a sample from the MIC benchmark, where $x_{S},x_{R}$ represents the moral situation and RoT, respectively. $y_{M}$ is its associated MF label. Each $y_{M}$ may contain between one and six MFs. $y_{J}$ is the moral judgment. We use $|y_{M}|=i$ $(1\leq i\leq 6,x\in\mathbb{Z})$ to indicate that $x$ is associated with $i$ MFs. We compute accuracy for each value of $i$.
Figure 1: Prompt Questions For the Task of Moral Foundations Classification. All content within { } is provided by the benchmark.
For baseline approaches, we follow prior work Forbes et al. (2020); Ziems et al. (2022) and consider two settings: (1) The base setting. For the moral judgment task, we compute $y_{J}=f_{\theta}(x_{S})$; and for MFC, we compute $y_{M}=f_{\theta}(x_{R})$. (2) The base+ setting. For the moral judgment task, we compute $y_{J}=f_{\theta}(x_{S},y_{M})$; and for MFC task, we compute $y_{M}=f_{\theta}(x_{R},d_{m})$. We additionally include the definitions of the six MFs ($d_{m}$) in the prompts. When the objective is moral judgment, we also explicitly provide the ground-truth MFs in the prompts. In other words, we intentionally provide the models with all available information in order to build strong baselines, which are then used to compare with the results of our framework. The setting of our proposed method is in line with the base+ setting but we incorporate an additional inference process in the input, e.g., $y_{J}=f_{\theta}(x_{S},y_{M},\{\text{inference}\})$.
To examine how moral reasoning performance scales with the size of the training corpus, we consider three data settings. The benchmark reports agreement levels, and 26% of the samples receive full agreement between all annotators. We use subsets of 5K, 10K, and all of the 23500 samples of the full-agreement data. All performance results are reported based on evaluation using the development set. More details about hyperparameters and experimental setup are in Appendix A.1.
4 Pragmatic Inference
In this section, we introduce our pragmatic inference methods for different moral reasoning objectives. Please note, pragmatic inferences often have more than one pathway to achieve a conclusion. Our approach represents one such pathway. Specifically, we build three inferential methods for moral foundation classification, moral judgment based on ground-truth moral foundations, and the joint task of moral foundations classification and moral judgment, respectively.
4.1 Motivation
Proposing a pragmatic inference method for moral reasoning is challenging because there is no linguistic theory that clearly defines what constitutes an appropriate pragmatic inference for LLMs. Although several theories describe how humans make moral judgments, LLMs lack human-like linguistic competence Bender and Koller (2020); Mahowald et al. (2024), making these theories unsuitable for direct application. On the other hand, LLMs are well known to capture distributional knowledge (i.e., distributional semantics) through statistical associations, and prior works have shown that they cannot leverage this distributional knowledge for satisfactory moral reasoning performance Liu et al. (2025c); Lenci and Sahlgren (2023). However, the distributional knowledge is much beyond distributional semantics, this is because pre-training corpora also contains static and context-limited texts Graesser et al. (1994); Sap et al. (2022), forcing LLMs not only capture distributional semantics but also approximate pragmatic cues through shallow statistical correlations.
For instance, the statement “calling your uncle by his given name is rude” is acceptable in many East Asian cultures but may not be perceived the same way elsewhere. Because pre-training texts provide no explicit contextual cues about the cultural setting in which such statements are embedded, LLMs may learn this norm only by associating negative descriptors “is rude” with the action “calling your uncle by his name” even though they lack any actual cultural understanding Pawar et al. (2025). This implies that LLMs can leverage these superficial associations to predict certain pragmatic cues. However, the extent to which they can do so remains unknown, and we cannot rely on a purely semantics-driven approach to simplify the pragmatic inference. We hypothesize that MFT-driven pragmatic inference can help LLMs address the challenges outlined above, based on the following two perspectives.
Accuracy(#MFs=1) Accuracy(#MFs=2) Accuracy(#MFs=3) Average
Model Size base base+ ours base base+ ours base base+ ours base base+ ours
Llama-1B 5000 .501 .696 .890 .402 .545 .856 .396 .460 .806 .433 .567 .851
10000 .582 .593 .832 .489 .449 .778 .414 .365 .743 .495 .469 .784
23500 .552 .671 .770 .505 .514 .796 .387 .495 .743 .481 .560 .770
Llama-3B 5000 .537 .754 .851 .466 .653 .822 .423 .572 .798 .475 .659 .824
10000 .386 .558 .725 .336 .507 .649 .324 .451 .698 .349 .505 .691
23500 .653 .790 .839 .493 .645 .784 .451 .617 .743 .532 .684 .789
Mistral-7B 5000 .612 .649 .803 .474 .504 .634 .441 .469 .577 .509 .540 .671
10000 .456 .653 .754 .373 .466 .606 .401 .419 .536 .410 .513 .632
23500 .487 .669 .768 .442 .536 .604 .383 .446 .559 .437 .550 .643
Table 2: Moral Foundations Classification Performance Across Different Training Dataset Sizes. The majority of prompt-reply pairs involve no more than three moral foundations labels; therefore, we report the performance on cases containing fewer than four moral foundations. For example, #MFs=2 indicates performance on instances annotated with a label containing two moral foundations. The optimal performance is highlighted with a bold font.
From the linguistic standpoint, our design of pragmatic inference methods is guided by three principles: (1) MFs are common ground which underlie all interactions. This has been supported by the theory and experiments of MFs Graham et al. (2013); Abdulhai et al. (2024). (2) MFs are abstract values extracted from discursive situations and hence need to be reified as context-specific interpretations Graham et al. (2011b). It is these context-specific interpretations of MFs which provide effective semantic cues for the models to reason morally. (3) Moral judgment is the judgment of good or bad by its nature, more specifically, the good/bad evaluation placed on the outcome of an action Nozick (1968); Schaich Borg et al. (2006); Emelin et al. (2021). Therefore, it is important to build the connections between contexts, discursive interpretations of MFs, and the outcome (i.e., behavioral consequence).
From a statistical standpoint, Liu et al. (2025b) demonstrate that the pragmatic gap between moral situations and moral reasoning objectives introduces greater degradation in language modeling performance than purely semantics-driven tasks. This is because LLMs tend to develop spurious associations, between moral situations and reasoning objectives, that undermine distributional semantics. To address this, MFT provides a principled signal: all moral situations can be interpreted through its lens. The key, then, is to leverage our pragmatic inference methods to supply linguistic cues that appropriately map moral situations to the LLMs’ hidden state space that is characterized by MFs, thereby alleviating the degradation of distributional semantics.
Figure 2: Prompting Questions for the Task of Moral Judgment Based on Ground-truth Moral Foundations. All content within {} is provided by the benchmark.
4.2 Moral Foundations Classification (MFC)
Figure 1 presents the prompting questions of the inferential steps used in the MFC task. The input here is a simple RoT and the definitions of six MFs. The inference steps are built as such: (1) Step 1 identifies the action encoded by the RoT Forbes et al. (2020). This step prepares the target (action) for moral evaluation. (2) Step 2 elicits the outcome of the action, namely, the behavioral consequence. One element of pragmatic inference is social norms, the variable that is probably most important in this task, given that RoTs were provided here without context. Norms for moral evaluations feature the physical and emotional harm to others that an action may incur Huebner et al. (2010); Kádár and Márquez-Reiter (2015); Turiel (1983). Therefore, Step 2 asks for the behavioral consequence as the linguistic manifestation of the social norms. This step setting also aligns with prior NLP studies showing that outcomes help establish the link between an action and its moral judgment Lindner et al. (2017); Lindner and Möllney (2019). (3) Step 3 connects the action and its consequence to relevant MFs by textualizing their metapragmatic links.
Table 2 reports performance on development set instances with 1–3 moral foundation labels. Moral situations with more than three labels are relatively rare. We also report the average performance for each experimental setting. Across all training dataset sizes, model variants, and numbers of moral foundation labels, our method consistently outperforms the baseline methods. Comparing the results of base and base+, it is evident that incorporating MFs is beneficial, which aligns with findings from previous studies Ziems et al. (2022); Liu et al. (2025b).
Interestingly, for each model size, increasing the training dataset does not always yield performance gains. We believe one significant reason is that the additional data does not substantially expand the range of unique moral situations, and, as we discussed in the Section 1, moral reasoning depends on interpreting the implicatures of such situations Taguchi (2011, 2018). On the other hand, for each dataset size, using larger models also does not always result in improved performance for both the baseline and our method. We believe this is because a training dataset of the same size is insufficient to modify the knowledge that larger LLMs have already acquired during pre-training compared to smaller models. A notable feature of our proposed pragmatic inference is that, unlike the baseline methods whose performance drops significantly as the number of underlying MFs increases, our method exhibits much smaller performance degradation.
4.3 Moral Judgment
In this section, we first introduce our inference method for making moral judgments based on ground-truth moral foundations. We then present an alternative inference method for jointly performing moral foundation classification and moral judgment. This joint approach guides LLMs to automatically identify the underlying moral foundations and subsequently make moral judgments based on the identified foundations.
4.3.1 Moral Judgment Based on Ground-truth Moral Foundations
Figure 2 presents the prompting questions used in the second pragmatic inference method that we have designed for moral judgment based on ground-truth MFs. Different from the first MFC task, where RoTs are not contextualized, the input here consists of the prompt–reply pair along with the underlying MFs. In other words, we provide the specific context (i.e., prompt+reply) and examine whether LLMs can effectively make moral judgments.
The inference steps of the second task are: (a) Step 1-3 is similar to the inference steps in the first task (i.e., specifying actions and consequences), but the difference is that we only care about actions that are relevant to the given MFs in the second task. (b) Step 4 connects the conclusion to the relevant moral foundations by textualizing their metapragmatic links. (c) Step 5 identifies the sentiment toward the actions mentioned in the Reply. We consider this step to be crucial for signaling the moral judgment of the Reply for two reasons. First, MFT provides a lens for establishing a meta-pragmatic link between the Reply and the moral judgment by evaluating how the actions uphold or violate specific MFs. Second, the expressed sentiment reflects the Reply’s stance toward actions that obey or violate these MFs. For example, if the Reply exhibits a positive stance toward an action that violates certain MFs, we infer a moral judgment of disagreement. (d) Step 6 aggregates the evidence from the preceding inference steps and aligns it with the ground-truth moral judgment label.
Model Size base base+ ours
Llama-1B 5000 .151 .080 .708
10000 .498 .335 .728
23500 .179 .318 .710
Llama-3B 5000 .160 .304 .712
10000 .205 .252 .735
23500 .249 .390 .710
Mistral-7B 5000 .418 .366 .704
10000 .438 .434 .760
23500 .486 .358 .696
Table 3: Accuracy of the Moral Judgment Task with Ground-Truth Moral Foundations Across Training Dataset and Model Sizes. The optimal performance is highlighted with a bold font.
Table 3 presents the performance on the moral judgment task with ground-truth MFs. Again, our pragmatic inference method (ours) consistently outperforms both baseline methods (base, base+) across different training dataset sizes. Similar to the MFC task, increasing data size does not necessarily improve the model performance. Since moral judgment is a three-class classification task, where random guessing yields an accuracy of approximately 0.333, the baseline methods perform worse than random guessing in some settings of dataset and model sizes. In contrast, baseline methods perform much better than random guessing on the MFC task. The difference indicates that the pragmatic gap varies in size between the two tasks. This is understandable as the involvement of context (prompt+reply) in the second task brings about more variability in interpretations than the close alignment of RoT and normative interpretations, e.g., the RoT  “It is good to keep in touch with friends” can rather be easily attributed to the relational norms between friends. Nevertheless, the different pragmatic gap sizes give rise to a crucial question of how LLMs can infer both of them in one task.
4.3.2 Joint Moral Foundations Classification and Moral Judgment
Building on the promising results of the first two tasks, the third task teaches LLMs to jointly infer MFs from a given situation and make corresponding moral judgments (denoted as MFC-Judgment). Please note that in our first task, moral foundations classification, we classify MFs from the given RoT. For the prompting questions, we replace the step 3 and 4 in Figure 2 with the following prompt questions:
(3). Based on the consequences of these Actions, why are their underlying moral foundations classified as {moral_foundations} according to the provided DEFINITIONs?
(4). How do Actions influence (up-regulate or down-regulate) their underlying moral foundations based on the valence of their consequences—positive, negative, or neutral?
Therefore, we guide LLMs to learn to infer the underlying moral foundations, motivated by their strong performance on the MFC task (Figure 1). The input has taken one part each from the above two tasks, namely, the situation (Prompt+Reply) and the definitions of MFs. This is closer to the real-life setting where a speaker only hears the interaction and has some (implicit) knowledge of MFs.
Model Size Judgment MFC-Judgment
Llama-1B 5000 .708 .706
10000 .728 .718
23500 .710 .704
Llama-3B 5000 .712 .750
10000 .735 .748
23500 .710 .766
Mistral-7B 5000 .704 .694
10000 .760 .720
23500 .696 .704
Table 4: Accuracy Performance of the Moral Judgment Task and the Joint Task Across Different Training Dataset Sizes and Model Architectures. The moral reasoning objective of these two tasks is the same. The optimal performance is highlighted with a bold font.
MFC Judgment
Accuracy(#MFs=1) Accuracy(#MFs=2) Accuracy(#MFs=3) Average
Model Size CoT ours CoT ours CoT ours CoT ours CoT ours
Llama-1B 5000 .661 .890 .475 .856 .428 .806 .521 .851 .629 .706
10000 .677 .832 .545 .778 .468 .743 .564 .784 .652 .718
23500 .717 .770 .509 .796 .401 .743 .542 .770 .641 .704
Llama-3B 5000 .719 .851 .544 .822 .459 .798 .574 .824 .635 .750
10000 .700 .725 .507 .649 .441 .698 549 .691 .653 .748
23500 .727 .839 .533 .784 .414 .743 .558 .789 .665 .766
Mistral-7B 5000 .677 .803 .533 .634 .401 .577 .537 .671 .637 .694
10000 .653 .754 .515 .606 .523 .536 .563 .632 .628 .720
23500 .649 .768 .420 .604 .383 .559 .484 .643 .644 .704
Table 5: Performance on the Moral Foundations Classification and Moral Judgment using the CoT-driven versus our MFT-driven inference methods. The optimal performance is highlighted with a bold font. Please note, our reported performance on moral judgment (ours) is from the model trained through the joint task. Our MFT-driven approach outperforms CoT across all tasks, model sizes, and training data scales. The prompts used in these evaluations are shown in Appendix 7.
Table 4 reports the performance of the joint task in comparison to the moral judgment task where ground-truth MFs are provided. Given the same moral reasoning objective, i.e., making a moral judgment, LLMs achieve performance on the joint task that is competitive with the standalone moral judgment task. Notably, for the LLama-3B model, performance on the joint task surpasses that of the moral judgment task based on ground-truth MFs. Furthermore, performance on the joint task consistently exceeds that of the baseline methods reported in Table 3. These results demonstrate that the current pragmatic inference framework enables LLMs to automatically infer underlying MFs and make moral judgments that are competitive with those based on ground-truth MFs, demonstrating the effectiveness of our proposed pragmatic inference framework.
In summary, this section first introduces two pragmatic inference frameworks: one for classifying MFs from a RoT and another for making moral judgments based on ground-truth MFs. Building on these insights, we further propose a third pragmatic inference framework that enables LLMs to automatically infer MFs and make corresponding moral judgments. The experimental results demonstrate the effectiveness of our proposed methods.
5 Analysis
In the previous sections, we demonstrated the effectiveness of our proposed pragmatic inference methods. In this section, we present additional analyses to further characterize these methods and highlight their benefits.
MFC Judgment
#MFs=1 #MFs=2 #MFs=3 Average
.694 .274 .147 .579 .466
Table 6: Performance of Direct LLM Prompting. The prompts used in these evaluations are provided in Appendix 6.
5.1 Performance of Direct LLM Prompting
Table 6 presents the results obtained by directly prompting the advanced reasoning model§§§deepseek-reasoning on the two tasks. Notably, its MFC performance is even lower than that of the Llama-3B baseline reported in Table 2, and its judgment performance falls below all baseline methods reported in Table 3. These results show that moral reasoning remains a challenging task even for advanced LLMs.
5.2 Effectiveness of Inference Driven by Moral Foundations Theory
Recall that in Section 4.1, we motivated the use of MFT to guide the inference process from two perspectives: linguistics and statistics. To empirically evaluate the effectiveness of MFT in our inference methods, we conduct additional experiments that generate training data by prompting off-the-shelf LLMs in a standard Chain-of-Thought (CoT) setting. We then compare the moral foundations classification and the moral judgment performance of models trained with our proposed MFT-driven inference methods against those trained with CoT-based inference.
Table 5 reports the experimental results comparing CoT-driven inference with our proposed MFT-driven approach. Across all model sizes, training data scales, and both tasks, the MFT-driven inference consistently outperforms the CoT baseline. These findings highlight the effectiveness of incorporating MFT to guide the inference process.
5.3 Language Modeling
Figure 3: Perplexity of Fine-tuned Models Under Different Moral Reasoning Approaches on the Moral Foundations Classification (MFC) and Moral Judgment Tasks. Our proposed methods result in a much smaller perplexity than baseline methods, suggesting a better language modeling capability.
Although fine-tuning can introduce forgetting in LLMs Kirkpatrick et al. (2017); Liu et al. (2024) and degrade their general capabilities, Liu et al. (2025b) emphasize that fine-tuning for moral reasoning, when not accounting for the pragmatic nature of morality, tends to yield higher perplexity. This is also verified in other morality-relevant tasks such as social stereotype mitigation Liu et al. (2025a). To further demonstrate the benefits of our proposed pragmatic inference methods, we calculate the Perplexity of the optimal fine-tuned checkpoints of baseline approaches and our proposed approach. We use the WikiText dataset Merity et al. (2016) to compute perplexity. Figure 3 reports perplexity results across different training dataset sizes and tasks. Models fine-tuned with our pragmatic inference methods consistently exhibit substantially lower perplexity compared to baseline approaches.
5.4 Moral Foundation-wise Analysis
Figure 4: Performance of MFC-Judgment Across Individual MFs. We report results only on test situations involving a single underlying moral foundation for more precise analysis. The black dashed line indicates the proportion of cases corresponding to each moral foundation in the training dataset. The X-axis orders the MFs in ascending order of their proportions.
Since MFs play an essential role in our pragmatic inference methods, we wonder whether and how the different distribution of them may affect the performance of models in the MFC-Judgment task. Figure 4 illustrates MFs-wise performance, highlighting a correlation between the number of situations of each MF in the training dataset and the eventual accuracy of moral judgment. Notably, care accounts for about 40% of the training situations, leading to higher model performance in test situations involving this foundation. In contrast, performance is much lower for test situations involving the sanctity foundation. These empirical results suggest a space for our methods to further improve model performance, if given sufficient and balanced annotation data.
6 Discussion
Pragmatic Inference for Moral Reasoning. Different from existing approaches that train models to map moral situations with or without moral foundations to moral judgment  Hendrycks et al. (2020); Forbes et al. (2020); Jiang et al. (2021); Ziems et al. (2022); Jiang et al. (2025), this paper proposes several possible bridges that effectively link the two. The bridge is built by textualizing the often unsaid variables (e.g., social norms, metapragmatic evaluations) in pragmatic inference, hence closing the gap between distributional semantics leveraged by LLMs and moral reasoning. Including the inferential steps is found to significantly increase the model performance. The findings are in line with previous studies that develop other pragmatic inferential methods for toxic language detection tasks  Sap et al. (2020); Zhou et al. (2023); Chen and Wang (2025). In fact, the concept of entextualizing pragmatic inferential process may be widely applicable to other non-logical reasoning tasks, such as the understanding of metaphors, humor, and sarcasm, which LLMs have thus far been found deficient in Barattieri di SanPietro et al. (2023); Ruis et al. (2023). Textualzing the inference processes for various pragmatic tasks also help to reveal the types of pragmatic knowledge that can be taught to LLMs and how they can be incorporated into model training.
Moral Foundations Theory for Moral Reasoning. We should note that it is not always uncontroversial to involve MFs in the pragmatic inference, because: (1) there is criticism that moral foundations lack neurological evidence for its role in moral reasoning and may oversimplify how humans perform moral reasoning Haidt and Joseph (2011); Braicovich (2022); and (2) they cannot capture cultural differences in moralities Graham et al. (2009, 2011a). While the latter still awaits for examinations with mutilingual datasets, our study, nevertheless, finds the effectiveness of MFs in machine learning Reinig et al. (2024), despite their drawbacks in human reasoning. Additionally, there is also neurological evidence linking human neural responses to moral foundations in the process of moral judgment Hopp et al. (2023). This leaves a further question to future studies as to whether there is a MFs-free inferential pathway that allows LLMs to sufficiently emulate or even precede humans in moral reasoning.
7 Conclusion
In this paper, we explore methods to enhance generalization in moral reasoning by addressing the gap between the distributional semantics captured by LLMs and the inherently pragmatic nature of morality. Grounded in Moral Foundations Theory, we first propose two pragmatic inference methods: one for classifying moral foundations from a given Rule of Thumb, and another for making moral judgments based on a given moral foundation. Building on these, we introduce a unified pragmatic inference method that enables LLMs to automatically infer moral foundations from a given moral situation and generate corresponding moral judgments. Our experimental results and analyses demonstrate that the proposed pragmatic inference methods significantly improve moral reasoning performance, reduce perplexity compared to baseline approaches, and avoid reliance on shallow heuristics.
Appendix A Appendix
A.1 Experimental Details
We apply supervised fine-tuning with a learning rate of 5e-5 and a batch size of 24. For mistral-7B, we leverage the LoRA setting with rank: 64, lora alpha: 16, lora dropout: 0.1. All experiments are conducted for 5 epochs using three random seeds (1, 2, and 3), and we report the optimal result across three random seeds.  Template for fine-tuning LLMs is: base: The rule-of-thumb judgment is {RoT}. The moral foundations underlying the rule-of-thumb are {moral foundations}. base+: {Definition of moral foundations} The rule-of-thumb judgment is {RoT}. The moral foundations underlying the rule-of-thumb are {moral foundations}. ours: {Definition of moral foundations} The rule-of-thumb judgment is {RoT}. ###Inference: {inference} the moral foundations underlying the rule-of-thumb are {moral foundations}.
Figure 5: Definition of Six Moral Foundations
Figure 6: Prompts for Direct LLMs Prompting
Figure 7: Prompts for Standard CoT Prompting